A different car accelerative from 0.12 m/s to 0. 52 meter per second that celebration of the car was 0.0 40 m/s squared the work done to accelerate the car was 0.48 joules

calculate the resultant force force needed to accelerate the car explain why the car has a maximum speed


1
Expert's answer
2022-03-10T10:17:55-0500

The work done by a force

"W=Fd=F\frac{v_f^2-v_i^2}{2a}"

Hence, the required force

"F=\frac{2aW}{v_f^2-v_i^2}"

"F=\frac{2*0.040*0.48}{0.52^2-0.12^2}=0.15\:\rm N"